26 February 2018

A Blustery Day in Eureka

On Monday February 26th, the skies were mostly cloudy throughout the day, only somewhat lessening in the evening. The temperature throughout the day hovered around -28 degrees Celsius at both the PEARL Ridge Lab and at the Eureka Weather Station, and there were approximately 5 hours and 20 minutes of sunlight. While these conditions weren’t great for some of the team, everyone still managed to make the most of the day.

Since the winds at the Ridge Lab were too strong to work on the roof, Kristof postponed troubleshooting the PEARL-GBS tracker. During the day he performed the resolution, polarization, and straylight tests on the PEARL-GBS. The instrument will make dark current measurements overnight, and normal measurements will resume tomorrow.

Last night due to the cloudy weather conditions, the DIAL did not operate. DIAL is fully operational and Ghazal and Alexey have it ready for tonight's measurements.

Because of high winds, it was not possible to operate the tracker dome and take solar measurements with the Bruker 125HR. Instead Sébastien did a second HCl cell test. Once completed, Erik put the KBr beamsplitter in the Bruker FTS and did a HBr cell test. Sébastien processed the HCl cell test, and found that the modulation efficiency is comparable to that obtained before the instrument was winterized. Solar measurements will start as soon as weather permits.

Xin did further snow sampling at two sites with the help of Paul and Pierre; one near the stream bed just behind 0PAL, and the other one on the sea ice about halfway between the weather station and the PEARL Ridge Lab. Xin found the snow density on sea ice to be much higher than expected, and it took him about an hour to finish a 1 cm vertical resolution sampling profile due to the struggle of dealing with this dense snow pack.

In addition to helping Xin, Paul joined Pierre and Alexey in testing a small drone for remote sensing capabilities.

Overnight on the 25th, it seemed cloudy enough for Emily to try a CRL sky calibration measurement with the aperture of the telescope covered by a depolarizer. Despite 11 hours of calibration, the clouds never gave signals quite high enough to be a satisfactory calibration, so this will have to be repeated another time. During the day of the 26th, Emily measured laser power and monitored laser beam shapes, did a half-hour lamp calibration, and then began regular lidar measurements again, which have so far been going on for 7 hours.

Finally SPS continued normal operations and analysis of preliminary data by Tom was found to look promising.

In the evening the team worked on individual projects and hoped that the weather would be more cooperative in the morning.

Cheers,

Paul Jeffery
[On Behalf of the 2018 Canadian ACE/OSIRIS Arctic Validation Campaign Team]
Back to Daily Reports
Back to Home